2lb/min of a certain gas whose R=767ft-lb/lb-R undergoes a steady flow isentropic process from 50 psia, 90ft/lb and 300ft/s until the pressure and velocity are doubled, and the volume is reduced by one third of its inital value. Determine a)H b)U c) Work


1
Expert's answer
2021-04-01T01:09:50-0400

Velocity is directly proportional to the Temperature of a system.

This means that when the velocity is doubled, the Absolute Temperature is doubled


P1/P2 = 1/2

V2/V1 = 1/3


P1/P2 = (V2/V1)γ

1/2 = (1/3)γ

γ = 0.631


"W = \dfrac{1}{γ -1}(p_1V_1 - p_2V_2)"


"W = \dfrac{1}{0.631- 1} (150V_2 - 50V_2)"


"W = -2.71(100V_2) = -271V_2"


"\u2206U = -W = 271V_2"